Mark deposited $9,000 into two saving accounts bearing simple interest. One of the accounts has an interest rate of 8% while the other rate is 7%. If the total interest earned after one year is $700, find the amount deposited into each of the accounts
Answer by macston(5194)

.
E=amount in 8% account; S=amount in 7% account
.
E+S=$9000
S=$9000-E Use this to substitute for S below.
.
0.08E+0.07S=$700 Sub for S.
0.08E+0.07($9000-E)=$700
0.08E+$630-0.07E=$700 Subtract %630 from each side.
0.01E=$70 Divide each side by 0.01.
E=$7000 ANSWER 1: He put $7000 in the 8% account.
.
S=$9000-E=$9000-$7000=$2000 ANSWER 2: He put $2000 in the 7% account.
.
CHECK:
0.08E+0.07S=$700
0.08($7000)+0.07($2000)=$700
$560+$140=$700
$700=$700